OpenStudy (anonymous):

The label on the car's antifreeze container claims to protect the car between -30 degree Celsius and 130 degree Celsius, to convert Celsios temperature to Fahrenheit temperature, the formula is C=5/9 (F-32), write and solve the inequality to determine the Fahrenheit temperature range at which this antifreeze protects the car.

OpenStudy (anonymous):

-30 C< T < 130 C becomes convert to F degrees by using F = (9/5) C + 32 e.g. C = 100 becomes (9/5)(100) + 32 = 212 F boiling point of water